OpenStudy (anonymous):

A tent in the shape of a triangular prism has a square base with a side of 8 feet and a height of 6 feet. What is the volume of the tent?

OpenStudy (anonymous):

Plug these numbers into the volume formula of a prism, volume=1/2*length*width*height. length=width=8 volume = 1/2*8*8*6 = 192 ft3

OpenStudy (anonymous):

thanks